Shift summary for Evening shift, June 1

Arun Tripathi (arun@physics.ucla.edu)
Tue, 1 Jun 1999 22:28:05 -0700 (PDT)


Shift Summary for Evening shift, 06/01/99
On Shift: Arun Tripathi
Valmiki Prasad

As usual, lot of other people were presnt too.

Hours of beam delivered at all intensities: 0.0
Hours of beam delivered at nominal intensity: 0.0
Hours of 99 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 0.0 hours

Accounting of lost time:

No beam this shift. Party time in the begining, and cryo problems in TeV
B1 later.

Activities:

15:31 Elizabeth fixes the REG WEST position bit problem by replacing the
bad cable between hardline station and latch station.
16:15 Ed modified ktevana to eliminate ptkick initialization by blklod
when running online. That fixes PTKICK problem.
19:27 Rick Coleman and Arun made an access to NM@, swapped the base from
ACC2 counter.
19:30 Elizabeth Fixed cabling error in REG/ABS position bits
19:45 Val and Arun make an access to check that the laser crate
repeater LV can be turned on and off from upstairs.
19:50 Rick Coleman tells us that by next monday, we would get our old
Epicure beam stop status NM2BS1-2 back.
Until then, we can look at the interlock chassis lights (RR207).
Look at the logbook XIX, P47 for details on how.
22:00 CsI channels 1957,1536,1813,117 show low light, and 2299,1216 show
missing HCC bits, no monitor.
22:15 Sydney and JasonH make an access to fix 2299 and 1216
2299 : had a loose connection and they swapped another board to
fix it
1216 : Same as 2299. Board was swapped.
23:30 Laplace and Kessler measure accidental trigger timing wrt RF
23:59 Shanahan P. fixes sporadic capid problem with channel 2360 by
re-adjusting L1 delay.
00:00 new shift takes over. We still await beam.....

=========================================================================
Runs ended this shift:

13182 : Not much useful data

Tapes brought to FCC this shift: 0